# Batch learning equals online learning in Bayesian supervised learning

### v3

@@ Abstract.
Using functoriality of probabilistic morphisms, we prove that sequential and batch Bayesian inversions coincide in supervised learning models with conditionally independent (possibly non-i.i.d.) data [Le2025]. This equivalence holds without domination or discreteness assumptions on sampling operators. We derive a recursive formula for posterior predictive distributions, which reduces to the Kalman filter in Gaussian process regression. For Polish label spaces $\mathcal{Y}$} and arbitrary input sets $\mathcal{X}$}, we characterize probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via projective systems, generalizing Orbanz [Orbanz2011]. We revisit MacEachern’s dependent Dirichlet processes (DDP) [MacEachern2000] using copula-based constructions [BJQ2012] and show how to compute posterior predictive distributions in universal Bayesian supervised models with DDP priors.
@@ 1. Introduction
In this paper, we study posterior distributions, posterior predictive distributions, and universal priors in Bayesian supervised learning. In particular, we prove that batch learning equals online learning in Bayesian supervised learning (Theorems 3.1, 4.1, 4.3).
The question of whether batch learning equals online learning in Bayesian learning has a notable history and importance in mathematical statistics and machine learning. The formal study of updating statistical conclusions one observation at a time (online learning) is known as sequential analysis. The mathematical groundwork for sequential analysis was laid by Abraham Wald [Wald1947]. The explicit formalization of sequential updating in a Bayesian context for estimation problems came to prominence with the development of state-space models and filtering theory [SS2023]. Rudolf E. Kálmán is arguably the most important figure in the practical application of this principle. The Kalman Filter, introduced around 1960, is a perfect example of online Bayesian learning. It uses the posterior from the previous time step as the prior for the current time step to recursively estimate the state of a linear dynamic system. While the Kalman filter is a specific algorithm for Gaussian models, its conceptual basis is precisely the equivalence of batch and online updating. The broader theoretical treatment of this idea in Bayesian statistics is often attributed to Dennis V. Lindley and Adrian F.M. Smith. Their work in the 1970s on Bayesian hierarchical models and the structure of Bayesian inference helped formalize and popularize these recursive computational structures. For instance, their 1972 paper [LS1972] is a landmark in this area. For Bayesian models with conjugate priors, the online approach is simply a recursive way of performing the same computation as the batch approach, breaking it down into smaller, manageable steps without any loss of information or change in the final inference. To the best of our knowledge, until now, the most general available theorem stating that batch Bayesian learning equals online Bayesian learning relies on the assumption of the classical Bayes’ theorem, which assumes either discrete data or sampling operators that are dominated Markov kernels, see, e.g., [SS2023, §3.3].
@@
Our paper is organized as follows. In Section 2, we recall the concept of probabilistic morphisms, their useful properties, and Bayesian learning models for supervised learning (Definition 2.1). In Section 3, using categorical properties of probabilistic morphisms, we prove that sequential Bayesian inversions in Bayesian supervised learning models is the same as batch Bayesian inversion (Theorem 3.1). Using a projective system, we also derive a formula for Bayesian inversions of a universal supervised learning model $({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ if $\mathcal{X}$} is a finite set (Theorem 3.4). We illustrate Theorems 3.1 and 3.4 by computing the posterior distributions of Dirichlet processes (Example 3.6). In Section 4, we prove recursive formulas for posterior predictive distributions (Theorems 4.1, 4.3) and illustrate Theorem 4.3 with Gaussian process regressions (Example 4.4). Section 5, assuming that $\mathcal{Y}$} is a Polish space and $\mathcal{X}$} is an arbitrary set, characterizes probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via a projective system, defined by finite subsets in $\mathcal{X}$} and a countable generating algebra of the $\sigma$-algebra of $\mathcal{Y}$} (Theorem 5.5). In Section 6, we illustrate Theorem 5.5 with MacEachern’s Dependent Dirichlet Processes (DDP) priors (Theorem 6.1) and indicate how to compute posterior predictive distributions of universal Bayesian supervised learning models with DDP priors. In the last Section 7, we discuss our results and the concept of predictive consistency in Bayesian supervised learning.
@@ 2. Preliminaries
- We denote by ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ the set of all measurable mappings from a measurable space $\mathcal{X}$} to a measurable space $\mathcal{Y}$}, and by ${\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ the set of all probabilistic morphisms from $\mathcal{X}$} to $\mathcal{Y}$}. We regard ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ as a subset of ${\bf Probm}({\mathcal{X}},{\mathcal{Y}})$, identifying ${\mathcal{Y}}\ni y$ with the Dirac measure $\delta_{y}\in{\mathcal{P}}({\mathcal{Y}})$. This is possible, since the Dirac map ${\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}),y\mapsto\delta_{y}$, is measurable [Lawvere62],[Giry82, Theorem 1].
- For $T_{i}\in{\bf Probm}({\mathcal{X}}_{i},{\mathcal{X}}_{i+1})$, $i=1,2$, we have [Chentsov72, Lemma 5.5]
(2.1) $\overline{T_{2}\circ T_{1}}=(T_{2})_{*}\circ\overline{T_{1}}.$
- We denote by ${\mathcal{Y}}^{\mathcal{X}}$ the set of all mappings from $\mathcal{X}$} to $\mathcal{Y}$}. If $\mathcal{Y}$} be measurable space, then ${\mathcal{Y}}^{\mathcal{X}}$ is a measurable space with the cylindrical $\sigma$-algebra denoted by $\Sigma_{cyl}({\mathcal{Y}}^{\mathcal{X}})$.
@@
- For any $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ the linear mapping
@@
is measurable [Lawvere62], [Giry82, Theorem 1]. If $\mathcal{X}$} and $\mathcal{Y}$} are Polish spaces, then ${\mathcal{P}}({\mathcal{X}})$ and ${\mathcal{P}}({\mathcal{Y}})$ are Polish spaces endowed with the weak*-topology $\tau_{w}$, and their $\sigma$-algebra $\Sigma_{w}$ is the Borel $\sigma$-algebra ${\mathcal{B}}(\tau_{w})$. Furthermore, $P_{*}T$ is a $(\tau_{w},\tau_{w})$-continuous map [Giry82, Theorem 1]. Moreover, for any $T_{1}\in{\bf Probm}({\mathcal{X}}_{1},{\mathcal{X}}_{2})$, $T_{2}\in{\bf Probm}({\mathcal{X}}_{2},{\mathcal{X}}_{3})$ we have [Lawvere62], [Chentsov72, Lemma 5.10],[Giry82, Theorem 1], [JLT21, Proposition 5]
@@
A Bayesian learning model for the supervised inference problem SBI consists of a quadruple $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$, where $\mu_{\Theta}\in{\mathcal{P}}(\Theta)$ and ${\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ is a measurable mapping.
(1) For any $X_{m}=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$, the Bayesian statistical model $(\Theta,\mu_{\Theta},\mathfrak{m}^{m}\circ E_{X_{m}}\circ{\mathbf{p}},{\mathcal{Y}}^{m})$ parameterizes sampling distributions of $Y_{m}=(y_{1},\ldots,y_{m})\in{\mathcal{Y}}^{m}$, where $y_{i}$ is a label of $x_{i}$, with the sampling operator ${\mathbf{p}}_{X_{m}}:=\mathfrak{m}^{m}\circ E_{X_{m}}\circ{\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{m})$.
(2) For a training sample $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, the posterior distribution $\mu_{\Theta|S_{n}}\in{\mathcal{P}}(\Theta)$ after seeing $S_{n}$ is the value ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}\big(\Pi_{\mathcal{Y}}(S_{n})\big)$ of a Bayesian inversion ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$.
@@
(1) A Bayesian inversion ${\mathbf{q}}_{X_{n}}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{X_{n}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$ is defined $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-a.s. uniquely in the weak sense.  [^1] See [Le2025, Theorem 2.9], for the precise definition. In this paper, we need only Formula (2.8) for a characterization of a Bayesian inversion. In Theorem 4.1 below, we shall show that, given inputs $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, for each $T_{m}\in{\mathcal{X}}^{m}$, the posterior predictive distribution
@@ 3. Bayesian inversions in Bayesian supervised learning models
In this Section, we prove two theorems (Theorems 3.1 and 3.4) for computing Bayesian inversions in supervised learning model. We illustrate these theorems with Example 3.6.

### v4

@@ Abstract.
In this paper we study Bayesian supervised learning models proposed by Lê in [Le2025]. We show the existence of Bayesian inversions on universal Bayesian supervised learning models $({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},\\ {\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ for arbitrary input space $\mathcal{X}$}, Souslin label space $\mathcal{Y}$}, and prior probability measure $\mu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$. Using functoriality of probabilistic morphisms, we prove that sequential and batch Bayesian inversions coincide in supervised learning models with conditionally independent (possibly non-i.i.d.) data [Le2025]. This equivalence holds without domination or discreteness assumptions on sampling operators. We derive a recursive formula for posterior predictive distributions, which reduces to the Kalman filter in Gaussian process regression. For Polish label spaces $\mathcal{Y}$} and arbitrary input sets $\mathcal{X}$}, we characterize probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via projective systems, generalizing Orbanz [Orbanz2011]. We revisit MacEachern’s dependent Dirichlet processes (DDP) [MacEachern2000] using copula-based constructions [BJQ2012] and show how to compute posterior predictive distributions in universal Bayesian supervised models with DDP priors.
@@ 1. Introduction
In this paper, we study the existence of Bayesian inversions, properties of posterior distributions, posterior predictive distributions, and universal priors in Bayesian supervised learning. We show that if the label space $\mathcal{Y}$} is a Souslin space then for any input space $\mathcal{X}$} and any universal model for supervised learning $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ there exist Bayesian inversions defining the learning algorithm for the problem SBI in the sense of Definition 2.1 (Theorem 3.1). We prove that batch learning equals online learning in Bayesian supervised learning (Theorems 3.4, 4.1, 4.3).
The question of whether batch learning equals online learning in Bayesian learning has a notable history and importance in mathematical statistics and machine learning. The formal study of updating statistical conclusions one observation at a time (online learning) is known as sequential analysis. The mathematical groundwork for sequential analysis was laid by Abraham Wald [Wald1947]. The explicit formalization of sequential updating in a Bayesian context for estimation problems came to prominence with the development of state-space models and filtering theory [SS2023]. Rudolf E. Kálmán is arguably the most important figure in the practical application of this principle. The Kalman Filter, introduced around 1960, is a perfect example of online Bayesian learning. It uses the posterior from the previous time step as the prior for the current time step to recursively estimate the state of a linear dynamic system. While the Kalman filter is a specific algorithm for Gaussian models, its conceptual basis is precisely the equivalence of batch and online updating. The broader theoretical treatment of this idea in Bayesian statistics is often attributed to Dennis V. Lindley and Adrian F.M. Smith. Their work in the 1970s on Bayesian hierarchical models and the structure of Bayesian inference helped formalize and popularize these recursive computational structures. For instance, their 1972 paper [LS1972] is a landmark in this area. For Bayesian models with conjugate priors, the online approach is simply a recursive way of performing the same computation as the batch approach, breaking it down into smaller, manageable steps without any loss of information or change in the final inference. To the best of our knowledge, until now, the most general available theorem stating that batch Bayesian learning equals online Bayesian learning relies on the assumption of the classical Bayes’ theorem, which assumes either discrete data or sampling operators that are dominated Markov kernels, see, e.g., [SS2023, Section 3.3].
@@
Our paper is organized as follows. In Section 2, we recall the concept of probabilistic morphisms, their useful properties, and Bayesian learning models for supervised learning (Definition 2.1). In Section 3, using categorical properties of probabilistic morphisms, we prove that sequential Bayesian inversions in Bayesian supervised learning models is the same as batch Bayesian inversion (Theorem 3.4). Using a projective system, we also derive a formula for Bayesian inversions of a universal supervised learning model $({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ if $\mathcal{X}$} is a finite set (Theorem 3.7). We illustrate Theorems 3.4 and 3.7 by computing the posterior distributions of Dirichlet processes (Example 3.9). In Section 4, we prove recursive formulas for posterior predictive distributions (Theorems 4.1, 4.3) and illustrate Theorem 4.3 with Gaussian process regressions (Example 4.4). Section 5, assuming that $\mathcal{Y}$} is a Polish space and $\mathcal{X}$} is an arbitrary set, characterizes probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ via a projective system, defined by finite subsets in $\mathcal{X}$} and a countable generating algebra of the $\sigma$-algebra of $\mathcal{Y}$} (Theorem 5.5). In Section 6, we illustrate Theorem 5.5 with MacEachern’s Dependent Dirichlet Processes (DDP) priors (Theorem 6.1) and indicate how to compute posterior predictive distributions of universal Bayesian supervised learning models with DDP priors. In the last Section 7, we discuss our results and the concept of predictive consistency in Bayesian supervised learning.
@@ 2. Preliminaries
- We denote by ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ the set of all measurable mappings from a measurable space $\mathcal{X}$} to a measurable space $\mathcal{Y}$}, and by $\operatorname{Probm}({\mathcal{X}},{\mathcal{Y}})$ the set of all probabilistic morphisms from $\mathcal{X}$} to $\mathcal{Y}$}. We regard ${\bf Meas}({\mathcal{X}},{\mathcal{Y}})$ as a subset of $\operatorname{Probm}({\mathcal{X}},{\mathcal{Y}})$, identifying ${\mathcal{Y}}\ni y$ with the Dirac measure $\delta_{y}\in{\mathcal{P}}({\mathcal{Y}})$. This is possible, since the Dirac map ${\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}),y\mapsto\delta_{y}$, is measurable [Lawvere62],[Giry82, Theorem 1].
- We denote by ${\mathcal{Y}}^{\mathcal{X}}$ the set of all mappings from $\mathcal{X}$} to $\mathcal{Y}$}. If $\mathcal{Y}$} is a measurable space, then ${\mathcal{Y}}^{\mathcal{X}}$ is a measurable space with the cylindrical $\sigma$-algebra denoted by $\Sigma_{cyl}({\mathcal{Y}}^{\mathcal{X}})$.
@@
- For any $T\in\operatorname{Probm}({\mathcal{X}},{\mathcal{Y}})$ the linear mapping
@@
is measurable [Lawvere62], [Giry82, Theorem 1]. If $\mathcal{X}$} and $\mathcal{Y}$} are Polish spaces, then ${\mathcal{P}}({\mathcal{X}})$ and ${\mathcal{P}}({\mathcal{Y}})$ are Polish spaces endowed with the weak*-topology $\tau_{w}$, and their $\sigma$-algebra $\Sigma_{w}$ is the Borel $\sigma$-algebra ${\mathcal{B}}(\tau_{w})$. Furthermore, $P_{*}T$ is a $(\tau_{w},\tau_{w})$-continuous map [Giry82, Theorem 1]. Moreover, for any $T_{1}\in\operatorname{Probm}({\mathcal{X}}_{1},{\mathcal{X}}_{2})$, $T_{2}\in\operatorname{Probm}({\mathcal{X}}_{2},{\mathcal{X}}_{3})$ we have [Lawvere62], [Chentsov72, Lemma 5.10],[Giry82, Theorem 1], [JLT21, Proposition 5]
@@
- For $T_{i}\in\operatorname{Probm}({\mathcal{X}}_{i},{\mathcal{X}}_{i+1})$, $i=1,2$, we have [Chentsov72, Lemma 5.5]
(2.3) $\overline{T_{2}\circ T_{1}}=(T_{2})_{*}\circ\overline{T_{1}}.$
@@
A Bayesian learning model for the supervised inference problem SBI consists of a quadruple $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$, where $(\Theta,\mu_{\Theta})$ is a probability parameter space, and ${\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ is a measurable mapping.
(1) For $X_{m}=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$, the sampling operator ${\mathbf{p}}_{X_{m}}:=\mathfrak{m}^{m}\circ E_{X_{m}}\circ{\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{m})$ parameterizes the sampling distributions of $Y_{m}=(y_{1},\ldots,y_{m})\in{\mathcal{Y}}^{m}$, where $y_{i}$ is a label of $x_{i}$, with certainty encoded in $\mu_{\Theta}$.
(2) For a training sample $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, the posterior distribution $\mu_{\Theta|S_{n}}\in{\mathcal{P}}(\Theta)$ after seeing $S_{n}$ is the value ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}\big(\Pi_{\mathcal{Y}}(S_{n})\big)$ where ${\mathbf{q}}_{\Pi_{\mathcal{X}}(S_{n})}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ is a Bayesian inversion of the Markov kernel ${\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$.
@@
(1) A Bayesian inversion ${\mathbf{q}}_{X_{n}}:{\mathcal{Y}}^{n}\to{\mathcal{P}}(\Theta)$ of the Markov kernel ${\mathbf{p}}_{X_{n}}:\Theta\to{\mathcal{P}}({\mathcal{Y}}^{n})$ relative to $\mu_{\Theta}$ is defined $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-a.s. uniquely in weak sense.  [^1] See [Le2025, Theorem 2.9], for the precise definition. In this paper, we need only Formula (2.8) for a characterization of a Bayesian inversion. In Theorem 4.1 below, we shall show that, given inputs $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, for each $T_{m}\in{\mathcal{X}}^{m}$, the posterior predictive distribution
@@ 3. Bayesian inversions in Bayesian supervised learning models
In this Section, first we prove Theorem 3.1 on the existence of Bayesian inversions on universal Bayesian supervised learning models. Then we prove Theorems 3.4 and 3.7 for computing Bayesian inversions in supervised learning model. We illustrate these theorems with Example 3.9.
Theorem 3.1 (Existence of Bayesian inversions).
Assume that $({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\\ \operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ is a universal Bayesian supervised learning model where $\mathcal{Y}$} is a Souslin space. Then for any $m\in{\mathbb{N}}$ and $X_{m}\in{\mathcal{X}}^{m}$ there exists a Bayesian inversion ${\mathbf{q}}^{(m)}:{\mathcal{Y}}^{m}\to{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ of the Markov kernel $\mathfrak{m}^{m}\circ E_{X_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ relative to $\mu$.
Proof.
Recall that a measurable space $({\mathcal{X}},\Sigma_{\mathcal{X}})$ is called a Souslin measurable space, if $({\mathcal{X}},\Sigma_{\mathcal{X}})$ is Borel isomorphic to a measurable space $({\mathcal{Y}},{\mathcal{B}}({\mathcal{Y}}))$ where $\mathcal{Y}$} is a Souslin metrizable space. Assume the condition of Theorem 3.1. By Theorem due to Dellacherie-Meyer [DM1978, Theorem 68, p. 76-III], there exists a bimeasurable bijection $f:{\mathcal{Y}}\to{\mathcal{Y}}_{met}$, where ${\mathcal{Y}}_{met}$ is a metrizable Souslin space. In fact, we can assume that $\mathcal{Y}$} is Borel isomorphic to a Souslin subset of $\mathbb{R}$}, see, e.g., Bogachev [Bogachev2007, Theorem 6.7.4, p. 25]. By [DM1978, Theorem 60, p. 73-III], ${\mathcal{P}}({\mathcal{Y}})$ is a Souslin measurable space. Thus, for any finite subset $A\subset{\mathcal{X}}$ such that $X_{m}\in A^{m}$ we can apply Jost-Lê-Tran theorem, which we recall in Proposition 3.2 below, to find an explicit formula for Bayesian inversion ${\mathbf{q}}_{X_{m},A}:{\mathcal{Y}}^{m}\to{\mathcal{P}}({\mathcal{Y}})^{A}$ of the Markov kernel $\mathfrak{m}^{m}\circ E_{X_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ to ${\mathcal{P}}({\mathcal{Y}})^{A}$.
Proposition 3.2.
[JLT21, Theorem 1] Suppose that $\mathcal{X}$} is a subset of a connected complete finite dimensional Riemannian manifold $(M^{n},g)$ equipped with the metric space structure induced by $g$, and $\Theta$ is a Souslin measurable space with a given probability measure $\mu_{\Theta}$. Assume that ${\mathbf{p}}:\Theta\to{\mathcal{P}}({\mathcal{X}})$ is a Markov kernel. For any $k\in{\mathbb{N}}^{+}$ and ${\bf x}=(x_{1},\cdots,x_{k})\in{\mathcal{X}}^{k}$ let $D_{r}(x_{i})$ denote the open ball of radius $r$ centered at $x_{i}\in{\mathcal{X}}$. Let $\mu_{\mathcal{X}}\in{\mathcal{P}}({\mathcal{X}})$ be the marginal measure on $\mathcal{X}$}. Then there exists a measurable subset $S\subset{\mathcal{X}}$ of zero $\mu_{\mathcal{X}}$-measure independent of $\bf x$} and a family of posterior distributions $\mu_{\Theta|{\mathcal{X}}}(\cdot|{\bf x})$ on $\Theta$ after seeing data ${\bf x}\in{\mathcal{X}}^{k}$ such that
(3.1) $\mu_{\Theta|{\mathcal{X}}}(B|{\bf x})=\lim_{r\to 0}\frac{\int_{B}\Pi_{i=1}^{k}{\mathbf{p}}(D_{r}(x_{i})|\theta)d\mu_{\Theta}(\theta)}{\int_{\Theta}\Pi_{i=1}^{k}{\mathbf{p}}(D_{r}(x_{i})|\theta)d\mu_{\Theta}(\theta)}$
for any $B$ in the Borel $\sigma$-algebra ${\mathcal{B}}(\Theta)=\Sigma_{\Theta}$, and for any ${\bf x}\in({\mathcal{X}}\setminus S)^{k}$. For ${\bf x}\in{\mathcal{X}}^{k}\setminus({\mathcal{X}}\setminus S)^{k}$ we assign an arbitrary value for $\mu_{\Theta|{\mathcal{X}}}(\cdot|{\bf x})\in{\mathcal{P}}(\Theta)$.
To complete the proof of Theorem 3.1 we shall show that a system of Bayesian inversions $\{{\mathbf{q}}_{X_{m},A},A\in\mathrm{P_{fin}({\mathcal{X}})}\}$ generates the required Bayesian inversion of the Markov kernel $\mathfrak{m}^{m}\circ E_{X_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{m})$. For the simplicity of notation we shall prove Theorem 3.1 for the case $m=1$. The general case follows from Theorem 3.4 in the next Section (Online formula for Bayesian inversions), or the reader can repeat verbatim the proof for the case $m=1$ to the case $m\geq 2$.
For $x\in{\mathcal{X}}$ let $\mathrm{P_{fin}}({\mathcal{X}},x):=\{A\in\mathrm{P_{fin}}({\mathcal{X}}),x\in A\}$. Then $\mathrm{P_{fin}}({\mathcal{X}},x)$ is a directed subset of the directed set $(\mathrm{P_{fin}}({\mathcal{X}}),\leq)$. Denote by $R^{\mathcal{X}}_{A}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{A}$ the restriction map. Clearly, $R^{\mathcal{X}}_{A}$ is a measurable map. Since $A$ is a finite set, the set ${\mathcal{P}}({\mathcal{Y}})^{A}$ is a Souslin measurable space.
Lemma 3.3.
Let $x\in{\mathcal{X}}$. Assume that $A\in\mathrm{P_{fin}}({\mathcal{X}},x)$. Let ${\mathbf{q}}^{A}_{x}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}})^{A}$ be a Bayesian inversion of $E_{x}^{A}:{\mathcal{P}}({\mathcal{Y}})^{A}\to{\mathcal{P}}({\mathcal{Y}})$ relative to the restriction $(R^{\mathcal{X}}_{A})_{*}\mu$ of $\mu$ to ${\mathcal{P}}({\mathcal{Y}})^{A}$ that is defined by Equation (3.1), i.e., for any $y\in{\mathcal{Y}}$ and $B\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{A}}$ we have
(3.2) ${\mathbf{q}}^{A}_{x}(B|y)=\lim_{r\to 0}\frac{\int_{B}E_{x}^{A}\big(D_{r}(y)|\theta\big)d(R^{\mathcal{X}}_{A})_{*}\mu(\theta)}{\int_{{\mathcal{P}}({\mathcal{Y}})^{A}}E_{x}\big(D_{r}(y)|\theta\big)d(R^{\mathcal{X}}_{A})_{*}\mu(\theta)}$
where $D_{r}(y)$ denotes the open ball of radius $r$ centered at $y$ in $\mathcal{Y}$}.
1) If $A_{1}\leq A_{2}$ then we have
(3.3) ${\mathbf{q}}^{A_{1}}_{x}=(R^{A_{2}}_{A_{1}})_{*}\circ{\mathbf{q}}^{A_{2}}_{x}.$
2) There exists a measurable measurable map ${\mathbf{q}}_{x}:{\mathcal{Y}}\to{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for any $A\in\mathrm{P_{fin}}({\mathcal{X}},x)$ we have
(3.4) ${\mathbf{q}}^{A}_{x}=(R^{{\mathcal{X}}}_{A})_{*}\circ{\mathbf{q}}_{x}.$
Proof.
1) Assume $A_{1}\leq A_{2}$. From the identities
$R^{\mathcal{X}}_{A_{1}}=R^{A_{2}}_{A_{1}}\circ R^{\mathcal{X}}_{A_{2}},$
${\mathcal{P}}({\mathcal{Y}})^{A_{2}}=(R^{A_{2}}_{A_{1}})^{-1}\big({\mathcal{P}}({\mathcal{Y}})^{A_{1}}\big),$
we obtain the following equalities for any $y\in{\mathcal{Y}}$, $r>0$, and $B\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{A_{1}}}$:
$\int_{B}E_{x}^{A_{1}}\big(D_{r}(y)|\theta_{1}\big)d(R^{\mathcal{X}}_{A_{1}})_{*}\mu(\theta_{1})=\int_{(R^{A_{2}}_{A_{1}})^{-1}(B)}E_{x}^{A_{2}}\big(D_{r}(y)|\theta_{2}\big)d(R^{\mathcal{X}}_{A_{2}})_{*}\mu(\theta_{2}),$
$\int_{{\mathcal{P}}({\mathcal{Y}})^{A_{1}}}E_{x}^{A_{1}}\big(D_{r}(y)|\theta_{1}\big)d(R^{\mathcal{X}}_{A_{1}})_{*}\mu(\theta_{1})=\int_{({\mathcal{P}}({\mathcal{Y}})^{A_{2}}}E_{x}^{A_{2}}\big(D_{r}(y)|\theta_{2}\big)d(R^{\mathcal{X}}_{A_{2}})_{*}\mu(\theta_{2}).$
It follows that for any $y\in{\mathcal{Y}}$ and $B\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{A_{1}}}$ we have
$\lim_{r\to 0}\frac{\int_{B}E_{x}^{A_{1}}\big(D_{r}(y)|\theta_{1}\big)d(R^{\mathcal{X}}_{A_{1}})_{*}\mu(\theta_{1})}{\int_{{\mathcal{P}}({\mathcal{Y}})^{A_{1}}}E_{x}^{A_{1}}\big(D_{r}(y)|\theta_{1}\big)d(R^{\mathcal{X}}_{A_{1}})_{*}\mu(\theta_{1})}=\lim_{r\to 0}\frac{\int_{(R^{A_{2}}_{A_{1}})^{-1}(B)}E_{x}^{A_{2}}\big(D_{r}(y)|\theta_{2}\big)d(R^{\mathcal{X}}_{A_{2}})_{*}\mu(\theta_{2})}{\int_{({\mathcal{P}}({\mathcal{Y}})^{A_{2}}}E_{x}^{A_{2}}\big(D_{r}(y)|\theta_{2}\big)d(R^{\mathcal{X}}_{A_{2}})_{*}\mu(\theta_{2})}.$
Hence, we obtain
${\mathbf{q}}^{A_{1}}_{x}(B|y)=(R^{A_{2}}_{A_{1}})_{*}{\mathbf{q}}^{A_{2}}_{x}(B|y).$
This proves (3.3).
2) The second assertion follows from the first assertion, taking into account the Kolmogorov extension theorem, see, e.g., [Bogachev2007, Corollary 7.7.2, p. 96]. To show that the map ${\mathbf{q}}_{x}$ is measurable, it suffices to show that for any $S\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}}$ the composition of the evaluation map $I_{S}:{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})\to{\mathbb{R}},\nu\mapsto\nu(S),$ with ${\mathbf{q}}_{x}$ is measurable. Since the $\sigma$-algebra of ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ is cylindrical, we can assume that $S=(R^{\mathcal{X}}_{A})^{-1}(S_{A})$ where $S_{A}\in\Sigma_{{\mathcal{P}}({\mathcal{Y}})^{A}}$ and $A\in\mathrm{P_{fin}}({\mathcal{X}})\}$. Then we have
[…]
Since ${\mathbf{q}}^{A}_{x}$ are measurable maps, $I_{S_{A}}\circ{\mathbf{q}}^{A}_{x}=I_{(R^{\mathcal{X}}_{A})^{-1}(S_{A})}\circ{\mathbf{q}}_{x}$ is measurable. This completes the proof of Lemma 3.3. ∎
[…]
(3.5) $(\Gamma_{\underline{{\mathbf{q}}_{x}}})_{*}\underline{E_{x}}_{*}\mu(B\times{\mathcal{A}})=(\Gamma_{\underline{E_{x}}})_{*}\mu({\mathcal{A}}\times B).$
Note that the LHS of (3.5) is equal to $(\Gamma_{\underline{{\mathbf{q}}^{A}_{x}}})_{*}\underline{E_{x}}_{*}\mu(B\times{\mathcal{A}})$. Since ${\mathbf{q}}^{A}_{x}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{A})$ is a Bayesian inversion of $E_{x}:{\mathcal{P}}({\mathcal{Y}})^{A}\to{\mathcal{P}}({\mathcal{Y}})$ , equality (3.5) holds. This completes the proof of Theorem 3.1.